OpenStudy (anonymous):

lim[x:1] ((√(x^(2)+1)-√(2))/(x-1))

myininaya (myininaya):

\[\lim_{x \rightarrow 1} \frac{\sqrt{x^2+1}-\sqrt{2}}{x-1}\]

myininaya (myininaya):

is that right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

did you try rationalizing the top?

OpenStudy (anonymous):

i got to lim\[\lim_{x \rightarrow 1} ((x+1)/(\sqrt{x ^{2}}+1) + \sqrt{2})\]

OpenStudy (anonymous):

but i dont know why the answer is \[2/2*\sqrt{2}\]

OpenStudy (anonymous):

ooohh! OKAY! thanks!

myininaya (myininaya):

what is x+x

myininaya (myininaya):

nvm you got it?

OpenStudy (anonymous):

yes! thanks!

OpenStudy (anonymous):

how do i give you a medal?